OpenStudy (winsteria):

FAN AND MEDAL The scatterplot and trend line show a positive correlation between advertising costs and sales. The slope was found using the points (6.4, 117) and (6.6, 120). What is the value of b in the equation of the trend line? y = 15x + b

OpenStudy (pawanyadav):

120=15×6.6+b You can calculate b from here So b=120-99=21

OpenStudy (winsteria):

Yes

OpenStudy (winsteria):

Thank you

OpenStudy (pawanyadav):

Don't mention it

OpenStudy (winsteria):

Could you help with this part of the question? I think that It goes with this one but I'm not sure.

OpenStudy (winsteria):

Given the equation of a trend line, can you find the exact amount of sales given the advertising cost?

OpenStudy (winsteria):

@Pawanyadav

OpenStudy (pawanyadav):

I can't understand what are you asking exactly You got the equation. y=15x+21 What after that?

OpenStudy (pawanyadav):

Just put y=1000 and find x

OpenStudy (winsteria):

How do I do that?

OpenStudy (pawanyadav):

What is the first equation

OpenStudy (winsteria):

y=15x+21

OpenStudy (winsteria):

y = 15x + b

OpenStudy (pawanyadav):

Is this question based on this equation If yes than Advertising cost is represented by x or y If it is x put x=1000 in the equation If it is y then put y=1000 in the equation and find the remaining one

OpenStudy (winsteria):

It Is

OpenStudy (pawanyadav):

Pls attach it

OpenStudy (winsteria):

@Pawanyadav

OpenStudy (winsteria):

Here

OpenStudy (pawanyadav):

Oh x is representing adv. cost and its 1000 So y=15000+21=15021